A diffraction grating, ruled with 300 lines per mm, is illuminated with a white light source at normal incidence.



(i) What is the angular separation, in the third-order spectrum, between the 400 nm and 600 nm lines? [5]




(ii) Water (of refractive index 1.33) now fills the whole space between the grating and the screen?


What is the angular separation, in the first-order spectrum, between the 400 nm and 600 nm lines? [5]


1
Expert's answer
2021-10-24T18:24:40-0400

1)

"d\sin \theta=k\lambda,"

"\sin\theta=\frac{k\lambda}d=0.45,"

"\theta=26.7°,"

2)

"d\frac{\sin \theta'}n=m\lambda,"

"\sin\theta'=\frac{mn\lambda}d=0.2,"

"\theta'=11.5°."